OpenStudy (anonymous):

Calculate the circumference of the circles with the following dimensions. C = circumference D = diameter R = radius PI = 3.1 R = 2 ft R = 14 ft R = 12 f

OpenStudy (mysesshou):

Similar to the other one :) use C=2*pi*R C=2 * 3.14 * 2 = C=2 * 3.14 * 14 = C=2 * 3.14 * 12 =

OpenStudy (mysesshou):

just plug in the radius values for R in the equation. :)